OpenStudy (anonymous):

5 green balls and 2 red balls together weigh 10 pounds, and 1 green ball and 4 red balls together weigh 7 pounds. If all red balls weigh the same amount and all green balls weigh the same, then what is the weight of 8 red and 8 green balls together? 5 green balls and 2 red balls together weigh 10 pounds, and 1 green ball and 4 red balls together weigh 7 pounds. If all red balls weigh the same amount and all green balls weigh the same, then what is the weight of 8 red and 8 green balls together? @Mathematics

OpenStudy (anonymous):

Let green ball's weight be x Let red ball's weight be y 5x + 2y = 10 x + 4y = 7 5x + 2y = 10 10x +4y = 20 - x +4y =7 9x = 13 x =13/9 y = 50/36 8 (13/9 + 50/36) = 68/3

OpenStudy (across):

I got something like 22.666...

OpenStudy (anonymous):

same 68/3 = 22.666...